The cweb Class

(Internal Interface)

Joachim Schrod^{∗}

August 29, 1995 (Revision 3.6 of cweb.cls)

1. This document shall describe the internal interface for TEX programmers who want to write packages or subclasses to enhance the cweb class. In particular, it shall describe all control sequences a programmer may redefine to adapt this class to special needs. (There are an awful lot of them, all cseqs that start with ‘Cweb’.)

Due to time constraints this description has not been written yet. It depends on the demand when it will be written. I.e., if you are a package/class writer and if you want to enhance the cweb class, send me email. Please mention the kind of enhancement you want to do. I will then add information to this description as needed.

Revision Log for cweb-conf.tex

Revision 2.1 (created at August 29, 1995 by Joachim Schrod)

Added (hint to non-existence of) description of the internal interface.

∗Email: schrod@iti.informatik.th-darmstadt.de